OpenStudy (anonymous):

log(4)(log(2)16) WILL MEDAL AND FAN PLEASEEE HELP

OpenStudy (anonymous):

\[\log_{4}(\log_{2}16) \]

OpenStudy (anonymous):

Also this problem if you can help me on this one too. its my last homework problems that i have no idea how to do

OpenStudy (campbell_st):

well 16 = 2^4 so for base 2 logs its log(16) = log(2^4) = 4log(2) = 4 because the base of the log and base of the exponential are the same, the value is 1. so then its base 4 logs for log(4) hopefullt this helps

OpenStudy (anonymous):

im even more confused now

OpenStudy (knowledge):

lets do this step by step what is \[\log _{2}16\]

OpenStudy (campbell_st):

srtart with the base 2 log rewrite 16 as 2 to the power 4 and apply the log law for power \[\log_{2}(16) = \log_{2}(2^4) = 4\log_{2}(2)\] and when the base of the log and the base of the exponent are the same the answer is 1 \[\log_{a}(a) = 1\] so \[\log_{2}(2) = 1\] \[4 \times \log_{2}(2) = 4 \times 1 = 4\] so you now have \[\log_{4}(4) = ?\] apply the same log law... base of the log and base of the exponent are the same...